The intensity of light varies inversely a the square of the distance from the light source. If the intensity from a light source 5 feet away is 8 lumens, what is the intensity at a distance of 2 feet?
Answer by stanbon(75887)

The intensity of light varies inversely a the square of the distance from the light source. If the intensity from a light source 5 feet away is 8 lumens, what is the intensity at a distance of 2 feet?
---------------------
I = k/d^2
----
Solve for "k":
8 = k/5^2
k = 200
-----
Equation:
I = 200/d^2
I = 200/2^2 = 200/4 = 50 lumens
--------------
Cheers,
Stan H.
--------------